A railway train 200m long passes over a bridge 600m long with a uniform speed of 36km/h. The time taken is
 
 

Grade:9

1 Answers

Harshit Singh
2 years ago
Dear Student

Speed in m/s = 36 x 5/18 = 10m/s

Distance to be covered = 800m
Time = distance/speed= 800/10 seconds
= 80 seconds or 1minute 20 sec

Thanks